OpenStudy (anonymous):

Write an equation for a circle with center at (2,3) and radius of 6 helpppppppp............

OpenStudy (anonymous):

\[(x-2)^2+(y-3)^2=36\]

OpenStudy (anonymous):

Circles have equations of the form \((x-h)^2+(y-k)^2=r^2\) where \((h,k)\) is the center and \(r\) is the radius.